OpenStudy (moazzam07):

What is the first step needed to solve 4 over 7 multiplied by x minus 5 equals negative 13? Subtract 13 from both sides Divide both sides by 7 Add 5 to both sides Multiply both sides by 4

OpenStudy (moazzam07):

please help

OpenStudy (moazzam07):

4 over 7 multiplied by x minus 5 equals negative 13?

OpenStudy (sdfgsdfgs):

what do u think should be the first step to solve \[\frac{ 4}{ 7}x-5=-13\]

OpenStudy (moazzam07):

\[\left(\begin{matrix}4 \\ 7\end{matrix}\right)x - 5 = -13 \]

OpenStudy (sdfgsdfgs):

so what u think should be the first step @moazzam07 ? :)

OpenStudy (moazzam07):

um add 5 to both sides

OpenStudy (sdfgsdfgs):

Good job @moazzam07 :)

OpenStudy (moazzam07):

thanks mate

OpenStudy (sdfgsdfgs):

welcome